Consider the demand for a good. At price Rs 4, the demand for the good is 25 units.


Suppose price of the good increases to Rs 5, and as a result, the demand for the good


falls to 20 units. Calculate the price elasticity?

1
Expert's answer
2021-11-18T10:56:04-0500

Price elasticity of demand (Ed​)=(−)Q
P​×△P
△Q​Here,P=Rs.4; P1​=Rs.5;△P=P1​−P=Rs.5−Rs.4=Rs.1 Q=25 units ; Q1​ = 20 units ; $$△Q=Q1​−Q=(20−25) units = (−)5unitsEd​=(−)25
4​×1
−5​=0.8.